In the quark model of fundamental particles, a proton is composed of three quarks: two "up" quarks, each having charge +2e/3, and one "down" quark, having charge -e/3. Suppose that the three quarks are equidistant from one another. Take the distance to be 1.20×10-15 m and calculate the potential energy of the subsystem of two "up" quarks. (MeV)
Answer #1

The potential energy of a system formed by two charge is

U=K\frac{q_{1}q_{2}}{r_{12}}

where r_{12}  is the distance between the charges. Therefore, if

q_{1}=q_{2}=\tfrac{2}{3}e=\mathrm{\tfrac{2}{3}\left( 1.6\times10^{-19}\,C \right)=1.07\times10^{-19}\,C}

the potential energy is

U=K\frac{q_{1}q_{1}}{r_{12}}

U=K\frac{q_{1}^{2}}{r_{12}}=K\frac{q_{2}^{2}}{r_{12}}

U=\mathrm{\left( 9\times10^{9}\,\tfrac{N.m^{2}}{C^{2}} \right )\frac{\left( 1.07\times10^{-19}\,C \right)^{2}}{\left( 1.20\times10^{-15}\,m \right)}}

U=\mathrm{8.59\times10^{-14}\,J}

but,

\mathrm{1\,eV=1.6\times10^{-19}\,J} \begin{matrix} && --> && \end{matrix} \mathrm{1\,J=\frac{1}{1.6\times10^{-19}}\,eV}

finally

U=\mathrm{8.59\times10^{-14}\,J}

U=\mathrm{8.59\times10^{-14}\left( \frac{1}{1.6\times10^{-19}}\,eV \right)}

U=\mathrm{0.537\,MeV}